# VersaSlot: Efficient Fine-grained FPGA Sharing with Big.Little Slots and Live Migration in FPGA Cluster

### v2

@@ I Introduction
Field Programmable Gate Arrays (FPGAs) have rapidly gained popularity in edge and data center computing, offering efficient application acceleration with higher flexibility and lower power consumption [1, 2, 3, 4, 5]. Leading cloud providers [6, 7, 8] have integrated FPGAs into their platforms, allowing users to develop highly customized and on-demand applications with Quality of Service. Meanwhile, modern FPGAs have significantly enhanced their capabilities with advanced manufacturing processes and increased logic elements, able to handle more complex and diverse computational demands [9, 10, 11]. These improvements facilitate fine-grained FPGA sharing [12, 13, 14, 15] in multi-user and multi-task environments.
@@
Traditional FPGA virtualization and sharing techniques [7, 16, 17] allocate the entire FPGA to a single application and implement the time-multiplexing by performing a full fabric reconfiguration. This approach introduces significant context switch overhead and often leads to FPGA under-utilization, as applications may not require the full FPGA resources. Instead, Dynamic Partial Reconfiguration (DPR) [18] enables an FPGA to split the programmable fabric into multiple slots, which can be reconfigured independently to host arbitrary application logic at runtime. It provides the capability to perform Partial Reconfiguration (PR) on a slot while other slots continue running. Currently, DPR has emerged as the most efficient and flexible technique for spatio-temporal FPGA sharing and virtualization [13, 14, 15] in the data center.
@@
However, current DPR-based spatio-temporal FPGA sharing still contains numerous overlooked challenges, leading to significant application performance degradation. First, in the current FPGA Processing System (PS), the PR module, PCAP (Processor Configuration Access Port), is limited to serial bitstream loading. This serial access causes frequent mutual blocking when multiple applications require PR concurrently. As the number of applications sharing an FPGA increases, the PR contention intensifies, resulting in highly prolonged and unpredictable application response times. Meanwhile, the mutual PR blocking further severely disrupts applications’ original pipelines, leading application tasks with dependencies in various slots to wait for each other to complete, preventing further execution. Recent systems like Nimblock [15] and DML[14] overlook these problems, resulting in application performance that falls short of the expected pipeline outcomes. Moreover, these works [15, 14] used single-core task scheduling, in which PR operations further block the launching of application tasks, leading to severe task blocking problems.
@@
Second, the uniform size of FPGA slots in previous work [14, 15] imposes strict constraints on task partitioning. For example, when optimizing application tasks using Xilinx’s High-Level Synthesis (HLS) tools, resource consumption typically exhibits stepwise increases rather than linear growth, which can easily cause resource over-subscription and under-utilization within slots. While prior work [12] attempted to address the problem by dynamically adjusting slot sizes based on application resource demand, the solution needs regenerating bitstreams at runtime, incurring large runtime overhead.
@@
However, similar to ARM’s big.LITTLE cores [19], the heterogeneous Big.Little slot design introduces increased complexity for resource allocation and scheduling. As for resource allocation, we propose an adaptive algorithm that combines primary allocation and redistribution along with binding and rebinding. The algorithm enables applications to flexibly switch between using Big and Little slots at runtime while also maximizing the utilization of available slot resources. The complementary between Big and Little slots can help mitigate under-utilization and over-subscription problems. In the context of scheduling, the VersaSlot dynamically bundles 3-in-1 tasks at runtime and schedules the corresponding tasks to available slots. The 3-in-1 task bundling can enhance the flexibility in task partitioning of applications. Importantly, VersaSlot decouples the PR from scheduling logic and distributes them across two CPU cores for asynchronous execution, thus effectively eliminating the task execution blocking problem.
@@ II Background
Dynamic Partial Reconfiguration (DPR), also known as Dynamic Function eXchange (DFX) from AMD/Xilinx [20], is an advanced FPGA configuration technique that enables dynamic reconfiguration of specific regions of the FPGA without interrupting the overall system operation. Instead, in traditional FPGA configuration flows, any functional modification requires reloading the entire bitstream file, leading to system downtime and a full restart, which causes large runtime overhead. The DFX process relies on the Processor Configuration Access Port (PCAP) [21] to load the partial bitstream from memory to the corresponding section of the FPGA. As it is essential to confirm the successful loading of the partial bitstream, DFX inherently prevents the PCAP from loading a new partial bitstream until the current one is fully loaded and blocks the corresponding CPU [21]. Meanwhile, to avoid unpredictable impacts on the remaining FPGA circuitry during the loading phase, a dedicated IP, DFX Decouple, is employed to decouple the reconfigurable region from the rest of the circuit. Recently, the DFX has been extensively used in FPGA virtualization and sharing in the cloud [13, 22, 14, 15], but limited to serial partial bitstream loading.
@@ III-A VersaSlot System Framework
The VersaSlot system, aligned with the typical structure of FPGA sharing design [15, 14], consists of two main components: the embedded hypervisor within the Processing System (PS) and the FPGA fabric layout in the Programmable Logic (PL), as illustrated in Figure 1. The VersaSlot hypervisor in the PS is implemented on ARM cores via bare-metal programming to minimize the system control overhead. Its primary responsibilities include scheduling and resource allocation, managing the loading of partial bitstreams and application data, and coordinating cross-board switching. For FPGA boards without a dedicated CPU, the hypervisor can run on the host CPU and control the FPGA via the PCIe interface.
@@
The Big.Little architecture is a heterogeneous slot design for spatio-temporal FPGA sharing, coupling resource-intensive Big slots and standard-resource Little slots. Although Little slots can increase resource granularity for more applications to share an FPGA, they also bring severe mutual PR contention and task execution blocking problems due to inherent DPR limitation in serial bitstream loading and CPU blocking. For example, as shown in Figure 2, two applications (App-1 and App-2), each with 3 tasks (T1, T2, T3) and batch sizes of 3 (B1, B2, B3) and 2 (B1, B2), are processed on an FPGA with 4 Little slots, in which each application is allocated with 2 slots that forms the execution pipeline. In the Nimblock [15], after the App-1 completes the PR of its first task, the App-2 starts the PR of its first task, during which it blocks the bitstream loading of App-1’s second task (T2) although App-1’s first batch item of the first task (T1B1) has finished the execution. Subsequently, the PR of App-1’s second task (T2) also blocks the PR of App-2’s second task. Moreover, since Nimblock handles all operations and scheduling by a single CPU core, the PR process of App-2’s T1 also blocks the task execution of App-1’s T1B2, and the PR of App-1’s T2 blocks the task execution of App-2’s T1B2. As the two applications continue running, mutual PR contention and task execution blocking alternate, disrupting each application’s original pipeline and significantly increasing response times. The figure shows only 4 slots with 2 applications, but as the number of shared slots and applications increases, this effect becomes more severe.
@@ III-C Slot Allocation and On-board Scheduling
Similar to the ARM big.LITTLE cores [19], the heterogeneous Big.Little slot architecture introduces greater complexity in resource allocation and scheduling. We propose an efficient slot allocation and scheduling algorithm for the architecture. When an application $A_{i}$ enters the candidate list, it is first added to the waiting list $C_{wait}$ for slot allocation. Once the application acquires slot resources $R_{A_{i}}$ (the maximum number of slots to use), all of its tasks are pushed into the ready list $Q_{T}$, awaiting scheduling for execution in Big or Little slots.
@@ 25             $L_{left}=L_{left}-\delta$;
As detailed in Algorithm 1, slot allocation is based on available slots and system status, featuring the processes of primary allocation and redistribution along with binding and rebinding. In the primary allocation, the system first allocates slots for applications in the waiting list $C_{wait}$. It prioritizes allocating Big slots to applications that can bundle tasks, as Big slots can significantly mitigate PR contention. After the Big slot allocation, Little slots are assigned to applications based on the optimal slot count $O_{A_{i}}^{L}$. The optimal slot allocation, derived through integer linear programming (ILP) as in [14, 15], represents the most efficient slot configuration for pipeline execution within either Little or Big slots, whose value is usually lower than the task count. The primary allocation ensures applications execute promptly along their task pipelines. When additional Little slots are left, Versaslot performs the redistribution based on applications already assigned with Little slots, prioritizing those at the front of the runnable queue to obtain the maximum needed slots. This redistribution effectively avoids slot idling and maximizes overall slot utilization. Meanwhile, to prevent situations where Big slots are idle while Little slots are overburdened, we propose a rebinding method for load balancing. For applications assigned to Little slots but not yet executing, we unbind them for new slot allocation, returning them to a waiting state. This ensures that Big slots can receive new applications for execution. But applications bound to the big slots can only complete all their tasks in the Big slots to avoid Big slot blocking caused by task dependencies.
@@ 24             PR$(T_{A_{i}})$ to an idle Little slot (Async in PR server);
Algorithm 2 outlines the overall scheduling process, which primarily includes 3-in-1 task online bundling, task execution launching, and scheduling tasks to their designated slots for PR. For newly added tasks in the ready list $Q_{T}$ assigned with Big slots, the scheduler bundles them into serial or parallel 3-in-1 tasks at runtime and replaces them with the bundled tasks in the ready list for execution. Meanwhile, the scheduler updates task execution statuses, loads task data, and launches batch executions. In previous work [14, 15], this step was often blocked by subsequent PR processes, leading to the task execution blocking problem. However, VersaSlot employs dual-core scheduling, which decouples the PR process and scheduling logic onto two separate CPU cores. When the scheduler assigns a new task to its corresponding available slot for PR, it sends an asynchronous request to the PR server without waiting for PR completion. Meanwhile, the scheduler limits application slot usage to its allocation, ensuring efficient spatio-temporal FPGA sharing. In VersaSlot’s Only.Little system, the scheduler follows the preemption mechanism as in [15], to prevent long-running tasks from monopolizing the FPGA. But the preemption brings more PR operations. The Big.Little applies preemption only in Little slots since an application cannot occupy both Big and Little slots simultaneously. Meanwhile, the inherent design of the slot redistribution in Big.Little also avoids monopolization.
@@ III-D2 Performance Degradation Metric $D_{\text{switch}}$
Inspired by Schmitt Trigger [23], we propose a switch loop for seamless cross-board switching across multiple FPGAs. As shown in Figure 4, as the $D_{\text{switch}}$ increases and reaches the threshold $T_{1}$, the system triggers the switching from the Only.Little slots to the Big.Little slots, thereby alleviating the PR contention. Similarly, when the $D_{\text{switch}}$ decreases to the threshold $T_{2}$, it triggers the switching from the Big.Little FPGA back to Only.Little FPGA, allowing more future applications to share an FPGA. The two user-configurable thresholds adjust switching sensitivity, with a buffer zone to prevent frequent switching and ensure system stability. Meanwhile, when the $D_{\text{switch}}$ metric enters the buffer zone, the system anticipates the direction of change, pre-configures the potential FPGA configuration, and loads task bitstreams into SD storage in a new FPGA. The pre-warming design further ensures seamless live migration with low overhead.
@@ IV Experiment and Evaluation
The VersaSlot system was evaluated on an FPGA cluster with two latest Xilinx UltraScale+ FPGAs (ZCU216). We used the same benchmark as previous work [15] and partitioned applications into multiple tasks based on the optimal fit between slot resources and task resource usage after synthesis. The applications include 3D Rendering (3DR, 3 tasks), LeNet (6 tasks), Image Compression (IC, 6 tasks), AlexNet (AN, 6 tasks), and Optical Flow (OF, 9 tasks). The partitioning and task bitstream generation were performed automatically by a TCL script in Vivado 2024.1. We compared the VersaSlot system with four existing scheduling algorithms/systems, including the traditional exclusive temporal multiplexing (Baseline) [7, 16], First-come-first-served spatio-temporal sharing (FCFS), round-robin (RR) from [22], and the state-of-the-art spatio-temporal sharing system Nimblock [15]. To simulate different congestion conditions, we randomly generated application workloads (10 sequences, 20 apps/sequence) with random batch size (5-30) and arrival intervals, including Loose (5000ms), Standard (1500ms-2000ms), Stress (150ms-200ms), and Real-time (50ms) for the evaluation.

### v3

@@ I Introduction
Field Programmable Gate Arrays (FPGAs) have rapidly gained popularity in edge and data center computing, offering efficient application acceleration with higher flexibility and lower power consumption [chung2018serving, wang2022via, wu2023fasda, zeng2023dfgas, zhao2023sagraph]. Leading cloud providers [alibabacloudWhatFPGA, amazonAmazonInstances, microsoftMicrosoftAzure] have integrated FPGAs into their platforms, allowing users to develop highly customized and on-demand applications with Quality of Service. Meanwhile, modern FPGAs have significantly enhanced their capabilities with advanced manufacturing processes and increased logic elements, able to handle more complex and diverse computational demands [bobda2022the, damiani2022blastfunction, ringlein2021acase]. These improvements facilitate fine-grained FPGA sharing [ahmed2018amorphos, zha2020virtualizing, dhar2022dml, mandava2023nimblock] in multi-user and multi-task environments.
@@
Traditional FPGA virtualization and sharing techniques [amazonAmazonInstances, putnam2014reconfigurable, tarafdar2017enabling] allocate the entire FPGA to a single application and implement the time-multiplexing by performing a full fabric reconfiguration. This approach introduces significant context switch overhead and often leads to FPGA under-utilization, as applications may not require the full FPGA resources. Instead, Dynamic Partial Reconfiguration (DPR) [dpr2020document] enables an FPGA to split the programmable fabric into multiple slots, which can be reconfigured independently to host arbitrary application logic at runtime. It provides the capability to perform Partial Reconfiguration (PR) on a slot while other slots continue running. Currently, DPR has emerged as the most efficient and flexible technique for spatio-temporal FPGA sharing and virtualization [zha2020virtualizing, dhar2022dml, mandava2023nimblock] in the data center.
@@
However, current DPR-based spatio-temporal FPGA sharing still contains numerous overlooked challenges, leading to significant application performance degradation. First, in the current FPGA Processing System (PS), the PR module, PCAP (Processor Configuration Access Port), is limited to serial bitstream loading. This serial access causes frequent mutual blocking when multiple applications require PR concurrently. As the number of applications sharing an FPGA increases, the PR contention intensifies, resulting in highly prolonged and unpredictable application response times. Meanwhile, the mutual PR blocking further severely disrupts applications’ original pipelines, leading application tasks with dependencies in various slots to wait for each other to complete, preventing further execution. Recent systems like Nimblock [mandava2023nimblock] and DML[dhar2022dml] overlook these problems, resulting in application performance that falls short of the expected pipeline outcomes. Moreover, these works [mandava2023nimblock, dhar2022dml] used single-core task scheduling, in which PR operations further block the launching of application tasks, leading to severe task blocking problems.
@@
Second, the uniform size of FPGA slots in previous work [dhar2022dml, mandava2023nimblock] imposes strict constraints on task partitioning. For example, when optimizing application tasks using Xilinx’s High-Level Synthesis (HLS) tools, resource consumption typically exhibits stepwise increases rather than linear growth, which can easily cause resource over-subscription and under-utilization within slots. While prior work [ahmed2018amorphos] attempted to address the problem by dynamically adjusting slot sizes based on application resource demand, the solution needs regenerating bitstreams at runtime, incurring large runtime overhead.
@@
However, similar to ARM’s big.LITTLE cores [armDocumentationx2013], the heterogeneous Big.Little slot design introduces increased complexity for resource allocation and scheduling. As for resource allocation, we propose an adaptive algorithm that combines primary allocation and redistribution along with binding and rebinding. The algorithm enables applications to flexibly switch between using Big and Little slots at runtime while also maximizing the utilization of available slot resources. The complementary between Big and Little slots can help mitigate under-utilization and over-subscription problems. In the context of scheduling, the VersaSlot dynamically bundles 3-in-1 tasks at runtime and schedules the corresponding tasks to available slots. The 3-in-1 task bundling can enhance the flexibility in task partitioning of applications. Importantly, VersaSlot decouples the PR from scheduling logic and distributes them across two CPU cores for asynchronous execution, thus effectively eliminating the task execution blocking problem.
@@ II Background
Dynamic Partial Reconfiguration (DPR), also known as Dynamic Function eXchange (DFX) from AMD/Xilinx [amdTechnicalInformation], is an advanced FPGA configuration technique that enables dynamic reconfiguration of specific regions of the FPGA without interrupting the overall system operation. Instead, in traditional FPGA configuration flows, any functional modification requires reloading the entire bitstream file, leading to system downtime and a full restart, which causes large runtime overhead. The DFX process relies on the Processor Configuration Access Port (PCAP) [amdPcap] to load the partial bitstream from memory to the corresponding section of the FPGA. As it is essential to confirm the successful loading of the partial bitstream, DFX inherently prevents the PCAP from loading a new partial bitstream until the current one is fully loaded and blocks the corresponding CPU [amdPcap]. Meanwhile, to avoid unpredictable impacts on the remaining FPGA circuitry during the loading phase, a dedicated IP, DFX Decouple, is employed to decouple the reconfigurable region from the rest of the circuit. Recently, the DFX has been extensively used in FPGA virtualization and sharing in the cloud [zha2020virtualizing, dario2020doos, dhar2022dml, mandava2023nimblock], but limited to serial partial bitstream loading.
@@ III-A VersaSlot System Framework
The VersaSlot system, aligned with the typical structure of FPGA sharing design [mandava2023nimblock, dhar2022dml], consists of two main components: the embedded hypervisor within the Processing System (PS) and the FPGA fabric layout in the Programmable Logic (PL), as illustrated in Figure 1. The VersaSlot hypervisor in the PS is implemented on ARM cores via bare-metal programming to minimize the system control overhead. Its primary responsibilities include scheduling and resource allocation, managing the loading of partial bitstreams and application data, and coordinating cross-board switching. For FPGA boards without a dedicated CPU, the hypervisor can run on the host CPU and control the FPGA via the PCIe interface.
@@
The Big.Little architecture is a heterogeneous slot design for spatio-temporal FPGA sharing, coupling resource-intensive Big slots and standard-resource Little slots. Although Little slots can increase resource granularity for more applications to share an FPGA, they also bring severe mutual PR contention and task execution blocking problems due to inherent DPR limitation in serial bitstream loading and CPU blocking. For example, as shown in Figure 2, two applications (App-1 and App-2), each with 3 tasks (T1, T2, T3) and batch sizes of 3 (B1, B2, B3) and 2 (B1, B2), are processed on an FPGA with 4 Little slots, in which each application is allocated with 2 slots that forms the execution pipeline. In the Nimblock [mandava2023nimblock], after the App-1 completes the PR of its first task, the App-2 starts the PR of its first task, during which it blocks the bitstream loading of App-1’s second task (T2) although App-1’s first batch item of the first task (T1B1) has finished the execution. Subsequently, the PR of App-1’s second task (T2) also blocks the PR of App-2’s second task. Moreover, since Nimblock handles all operations and scheduling by a single CPU core, the PR process of App-2’s T1 also blocks the task execution of App-1’s T1B2, and the PR of App-1’s T2 blocks the task execution of App-2’s T1B2. As the two applications continue running, mutual PR contention and task execution blocking alternate, disrupting each application’s original pipeline and significantly increasing response times. The figure shows only 4 slots with 2 applications, but as the number of shared slots and applications increases, this effect becomes more severe.
@@ III-C Slot Allocation and On-board Scheduling
Similar to the ARM big.LITTLE cores [armDocumentationx2013], the heterogeneous Big.Little slot architecture introduces greater complexity in resource allocation and scheduling. We propose an efficient slot allocation and scheduling algorithm for the architecture. When an application $A_{i}$ enters the candidate list, it is first added to the waiting list $C_{wait}$ for slot allocation. Once the application acquires slot resources $R_{A_{i}}$ (the maximum number of slots to use), all of its tasks are pushed into the ready list $Q_{T}$, awaiting scheduling for execution in Big or Little slots.
@@ 25     $L_{left}=L_{left}-\delta$;
As detailed in Algorithm 1, slot allocation is based on available slots and system status, featuring the processes of primary allocation and redistribution along with binding and rebinding. In the primary allocation, the system first allocates slots for applications in the waiting list $C_{wait}$. It prioritizes allocating Big slots to applications that can bundle tasks, as Big slots can significantly mitigate PR contention. After the Big slot allocation, Little slots are assigned to applications based on the optimal slot count $O_{A_{i}}^{L}$. The optimal slot allocation, derived through integer linear programming (ILP) as in [dhar2022dml, mandava2023nimblock], represents the most efficient slot configuration for pipeline execution within either Little or Big slots, whose value is usually lower than the task count. The primary allocation ensures applications execute promptly along their task pipelines. When additional Little slots are left, Versaslot performs the redistribution based on applications already assigned with Little slots, prioritizing those at the front of the runnable queue to obtain the maximum needed slots. This redistribution effectively avoids slot idling and maximizes overall slot utilization. Meanwhile, to prevent situations where Big slots are idle while Little slots are overburdened, we propose a rebinding method for load balancing. For applications assigned to Little slots but not yet executing, we unbind them for new slot allocation, returning them to a waiting state. This ensures that Big slots can receive new applications for execution. But applications bound to the big slots can only complete all their tasks in the Big slots to avoid Big slot blocking caused by task dependencies.
@@ 17     $Q_{T}=Q_{T}\setminus T_{A_{i}}$; $isFinished(A_{i})$$\Rightarrow S=S\setminus A_{i}$;
Algorithm 2 outlines the overall scheduling process, which primarily includes 3-in-1 task online bundling, task execution launching, and scheduling tasks to their designated slots for PR. For newly added tasks in the ready list $Q_{T}$ assigned with Big slots, the scheduler bundles them into serial or parallel 3-in-1 tasks at runtime and replaces them with the bundled tasks in the ready list for execution. Meanwhile, the scheduler updates task execution statuses, loads task data, and launches batch executions. In previous work [dhar2022dml, mandava2023nimblock], this step was often blocked by subsequent PR processes, leading to the task execution blocking problem. However, VersaSlot employs dual-core scheduling, which decouples the PR process and scheduling logic onto two separate CPU cores. When the scheduler assigns a new task to its corresponding available slot for PR, it sends an asynchronous request to the PR server without waiting for PR completion. Meanwhile, the scheduler limits application slot usage to its allocation, ensuring efficient spatio-temporal FPGA sharing. In VersaSlot’s Only.Little system, the scheduler follows the preemption mechanism as in [mandava2023nimblock], to prevent long-running tasks from monopolizing the FPGA. But the preemption brings more PR operations. The Big.Little applies preemption only in Little slots since an application cannot occupy both Big and Little slots simultaneously. Meanwhile, the inherent design of the slot redistribution in Big.Little also avoids monopolization.
@@ III-D2 Performance Degradation Metric $D_{\text{switch}}$
Inspired by Schmitt Trigger [wikipediaSchmittTriggerWikipedia], we propose a switch loop for seamless cross-board switching across multiple FPGAs. As shown in Figure 4, as the $D_{\text{switch}}$ increases and reaches the threshold $T_{1}$, the system triggers the switching from the Only.Little slots to the Big.Little slots, thereby alleviating the PR contention. Similarly, when the $D_{\text{switch}}$ decreases to the threshold $T_{2}$, it triggers the switching from the Big.Little FPGA back to Only.Little FPGA, allowing more future applications to share an FPGA. The two user-configurable thresholds adjust switching sensitivity, with a buffer zone to prevent frequent switching and ensure system stability. Meanwhile, when the $D_{\text{switch}}$ metric enters the buffer zone, the system anticipates the direction of change, pre-configures the potential FPGA configuration, and loads task bitstreams into SD storage in a new FPGA. The pre-warming design further ensures seamless live migration with low overhead.
@@ IV Experiment and Evaluation
The VersaSlot system was evaluated on an FPGA cluster with two latest Xilinx UltraScale+ FPGAs (ZCU216). We used the same benchmark as previous work [mandava2023nimblock] and partitioned applications into multiple tasks based on the optimal fit between slot resources and task resource usage after synthesis. The applications include 3D Rendering (3DR, 3 tasks), LeNet (6 tasks), Image Compression (IC, 6 tasks), AlexNet (AN, 6 tasks), and Optical Flow (OF, 9 tasks). The partitioning and task bitstream generation were performed automatically by a TCL script in Vivado 2024.1. We compared the VersaSlot system with four existing scheduling algorithms/systems, including the traditional exclusive temporal multiplexing (Baseline) [amazonAmazonInstances, putnam2014reconfigurable], First-come-first-served spatio-temporal sharing (FCFS), round-robin (RR) from [dario2020doos], and the state-of-the-art spatio-temporal sharing system Nimblock [mandava2023nimblock]. To simulate different congestion conditions, we randomly generated application workloads (10 sequences, 20 apps/sequence) with random batch size (5-30) and arrival intervals, including Loose (5000ms), Standard (1500ms-2000ms), Stress (150ms-200ms), and Real-time (50ms) for the evaluation.